# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v3

@@ C.1 Details of All Task Combinations
  • Walker-stand&\&&walk. This task combination includes two tasks: Walker-stand and Walker-walk, which both require a two-leg robot to maintain its height above a certain threshold, and Walker-walk also hopes it to move faster than a certain speed, i.e., their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\end{split}$ (30)
    Therefore, for all states, the optimal action of Walker-walk is also optimal in Walker-stand and TDR here is 0.
@@
  • Cartpole-balance&\&&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively. The optimal actions of these tasks are both hope to balance the agent and thus TDR here is 0.
  • Acrobot-swingup&\&&sparse. This task combination includes two tasks: Acrobot-swingup and Acrobot-swingup__\__sparse, which both hope to swing up a double pendulum with dense and sparse rewards respectively. Also, the optimal actions of these tasks are both hoped to swingup the agent and thus TDR here is 0.
  • Quadruped-walk&\&&run. This task combination includes two tasks: Quadruped-walk and Quadruped-run, which both require a quadruped agent to move faster than a target speed. The target speed of the running task is larger than the walking task, i.e., their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{walk}}=\text{tolerance}(\text{speed},(0.5,\infty)),\\ &\mathcal{R}_{\text{run}}=\text{tolerance}(\text{speed},(5.0,\infty)).\end{split}$ (31)
    Therefore, TDR here is still 0 since optimal actions of Quadruped-run are also optimal in Quadruped-walk.
  • Reacher-easy&\&&hard. This task combination includes two tasks: Reacher-easy and Reacher-hard, which both hope to control a two-link planar reacher to prostrate the target sphere. The sizes of target spheres in these two tasks are different, i.e., their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{easy}}=\text{tolerance}(\text{distance},(0.0,0.05)),\\ &\mathcal{R}_{\text{hard}}=\text{tolerance}(\text{distance},(0.0,0.015)).\end{split}$ (32)
    Thus TDR here is not 0 but still small.
  • Hopper-stand&\&&hop. This task combination includes two tasks: Hopper-stand and Hopper-hop, which both hope the height of a one-legged hopper is larger than the target height. While Hopper-stand hopes the control force is 0, Hopper-hop hopes to control the robot to hop forward and its speed is larger than the target speed. Their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{speed},(2.0,\infty)),\\ \mathcal{R}_{\text{hop}}&=\text{tolerance}(\text{height},(0.6,2))\\ &*\text{tolerance}(\text{control},(0.0,0.0)).\end{split}$ (33)
    Since one agent can not hop when controlling the force as 0, TDR here is not 0 but not huge.
  • Walker-stand&\&&prostrate. This task combination includes two tasks: Walker-stand and Walker-prostrate. Walker-stand hopes the height of a two-leg robot to be larger than the target height. And we design the task Walker-prostrate, which hopes its height to be lower than another target height. Their reward functions can be roughly described as
    $\begin{split}&\mathcal{R}_{\text{stand}}=\text{tolerance}(\text{height},(1.2,\infty)),\\ &\mathcal{R}_{\text{prostrate}}=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (34)
    Thus TDR here is huge since the optimal Q functions of these two tasks are almost the opposite.
@@
  • Walker-walk&\&&prostrate. This task combination includes two tasks: Walker-walk and Walker-prostrate. Walker-walk hopes the height of the improved planar walker to be larger than a target height, while Walker-prostrate hopes it to be lower than a target height. Moreover, Walker-walk hopes the speed of the robot is larger than another target speed, i.e., their reward functions can be roughly described as
@@
    $\begin{split}\mathcal{R}_{\text{walk}}&=\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1.0,\infty)),\\ \mathcal{R}_{\text{prostrate}}&=\text{tolerance}(\text{height},(0.0,0.2)).\end{split}$ (35)
    Thus TDR here is also huge since the optimal Q functions of these two tasks are almost the opposite.
Figure 4: Average cumulative reward curves over different task combinations in DMC for CURL, SAC+AE, PlaNet, Dreamer, and TAD. The x-axes indicate the number of episodes interacting with the environment, and the y-axes indicate the performance, including average rewards with std.
@@
Tasks Hypothesis Walker-stand&\&&walk Cartpole-balance&\&&sparse Acrobot-swingup&\&&sparse Quadruped-walk&\&&run
CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 996.8 $\pm$ 0.8 11.0 $\pm$ 1.8 203.3 $\pm$ 69.9
SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 993.4 $\pm$ 0.6 11.7 $\pm$ 5.4 203.2 $\pm$ 70.2
PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 412.8 $\pm$ 65.9 12.2 $\pm$ 9.8 305.3 $\pm$ 206.6
Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 997.1 $\pm$ 1.1 159.9 $\pm$ 108.4 345.4 $\pm$ 218.3
TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 998.9 $\pm$ 0.3 217.3 $\pm$ 107.1 505.0 $\pm$ 119.1
Tasks Hypothesis Reacher-easy&\&&hard Hopper-stand&\&&hop Walker-stand&\&&prostrate Walker-walk&\&&prostrate
CURL $\Pi_{1}$ 552.5 $\pm$ 88.9 137.1 $\pm$ 9.2 496.3 $\pm$ 8.7 441.0 $\pm$ 7.7
SAC+AE $\Pi_{1}$ 669.5 $\pm$ 99.4 118.2 $\pm$ 83.1 504.4 $\pm$ 6.7 445.8 $\pm$ 2.6
PlaNet $\Pi_{2}$ 341.0 $\pm$ 228.3 25.9 $\pm$ 25.7 544.3 $\pm$ 68.6 445.4 $\pm$ 5.1
Dreamer $\Pi_{2}$ 506.1 $\pm$ 382.9 436.9 $\pm$ 18.9 510.0 $\pm$ 6.7 449.7 $\pm$ 12.5
TAD $\Pi_{3}$ 823.3 $\pm$ 237.4 503.9 $\pm$ 22.2 941.1 $\pm$ 4.0 913.7 $\pm$ 11.5
@@ C.2 Details of All Task Distributions
  • Cheetah_normal-_\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
  • Pendulum_normal-_\__angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
@@
    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and testing tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
  • Walker_normal-_\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. As for the hyper-parameters, we follow previous works Ha and Schmidhuber (2018); Hafner et al. (2019b, a) and select 2 as the action repeat for all experiments following Hafner et al. (2019a).
C.3 Full Experimental Results for TDR
In tasks where TDR is 0, like Walker-stand&\&&walk and Cartpole-balance&\&&sparse, the optimal performance of TAD and baselines are similar since these tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, as derived in Theorem 2. In these cases, TAD still converges faster and shows more robust performance. In Acrobot-swingup&\&&sparse and Quadruped-walk&\&&run, although TDR of these tasks is 0, TAD can surpass baselines since it can better distinguish different tasks. Finally, in the last four environments with non-zero TDR, TAD achieves a notable improvement compared with baselines. Especially, in Walker-stand&\&&prostrate and Walker-walk&\&&prostrate, where TDR is significantly huge, baselines perform poorly as these policies belong to $\Pi_{1},\Pi_{2}$ and can not differentiate different tasks, which also aligns with our Theorem 2. In contrast, TAD can distinguish between tasks by utilizing reward signals and improves performance conspicuously.
C.4 Visualization Results for Task Generalization
@@
Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for testing tasks), we plot the speed of the agent as a function of the timestep in Fig. 5. As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in supplementary materials.
@@ Appendix D Experimental Details for MuJoCo
We here introduce state-based control tasks (Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D) in detail, following the setting of previous meta RL works Finn et al. (2017); Rakelly et al. (2019).
@@
  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 testing tasks for experiments.
  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 testing tasks for experiments.
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. We select 1 as the action repeat for all experiments following.
Appendix E Ablation Study
E.1 Dynamic Generalization
E.2 Sparse Reward
@@
Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed. In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 7, with the increasing of SR, although the performance of TAD decrease since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
Appendix F Ethics Issues and Border Impact

### v4

@@ C.1 Details of All Task Combinations
  • Cartpole-balance&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively. The optimal actions of these tasks are both hoped to balance the agent and thus TDR here is 0.
@@
  • Walker-stand&walk&prostrate&flip. This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip. Walker-stand hopes the height of a two-leg robot to be larger than the target height. Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed. Walker-prostrate hopes the height of the robot to be lower than the target height. Finally, Walker-flip hopes the robot to stand and move forward to the target speed by executing a rapid twist and jump. In detail, their reward functions can be roughly described as
@@
    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (31)
    In this situation, for all states, the optimal actions of Walker-walk/stand and Walker-prostrate are opposite, and TDR here is huge.
  • Cheetah-run&run__\__back&flip&flip__\__back. This task combination includes four tasks: Cheetah-run, Cheetah-run__\__back, Cheetah-flip, and Cheetah-flip__\__back. Cheetah-run hopes to control a running planar biped to run forward within a target speed. Cheetah-run__\__back, differently, hopes to control the Cheetah robot to run backward within a target speed. Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump. Similarly, Cheetah-flip__\__back hopes to control the robot to move backward by flipping.
    $\begin{split}\mathcal{R}_{\text{Cheetah}\_\text{run}}=&\text{tolerance}(\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{run$\_$back}}=&\text{tolerance}(-\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip}}=&\text{tolerance}(\text{angmomentum},(5,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip$\_$back}}=&\text{tolerance}(-\text{angmomentum},(5,\infty)).\end{split}$ (32)
    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run__\__back are opposite, and TDR here is also huge.
@@
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
Table 4: Performance (mean $\pm$ std) in DMC of CURL, CURL (w/ r), SAC+AE, and SAC+AE (w/ r).
Train&Test Train&Test Train Test Train Test
@@ C.2 Details of All Task Distributions
  • Cheetah__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
  • Pendulum__\__angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
@@
    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and test tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
  • Walker__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 3 days. As for the hyper-parameters, we follow previous works (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) and select 2 as the action repeat for all experiments following (Hafner et al. 2019a).
C.3 Visualization Results for Task Generalization
Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig. 6. As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in supplementary materials.
@@ Appendix D Experimental Details for MuJoCo
We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works (Finn, Abbeel, and Levine 2017; Rakelly et al. 2019).
@@
  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 test tasks for experiments.
  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 test tasks for experiments.
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 1 day. We select 1 as the action repeat for all experiments following.
@@
Appendix E Ablation Study
E.1 Ablation Study on Reward Signals
In this part, we do ablation studies on reward signals for model-free methods including CURL and SAC+AE, which both belong to $\Pi_{1}$. We design CURL (w/ r) and SAC+AE (w/ r) by directly adding reward signals into the observation based on CURL and SAC+AE, respectively. As shown in Table. 4, the performance of CURL (w/ r) and SAC+AE (w/ r) is similar to CURL and SAC+AE. The major reason is that CURL (w/ r) and SAC+AE (w/ r)only utilize the observation and reward of the current timestep without historical information. Thus it is still difficult for them to distinguish different tasks.
E.2 Dynamic Generalization
E.3 Sparse Reward
@@
Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed. In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 7, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
Appendix F Ethics Issues and Broader Impact